Question 9 (1 point) Find the percent of z-scores between z = - 2.06 and z = -1.23.
Answer #1

# percentage of Z-score is Between -2.06 and -1.23

ie P(-2.06<Z<-1.23)

P(-2.06<Z<-1.23)=P(Z<-1.23)-P(Z<-2.06)

=0.1093-0.0197=0.0896

#Value of z is obtain from standard normal table

P(-2.06<Z<-1.23) =0.0896 ie 8.96 %

  

#percentage of Z-score is Between -2.06 and -1.23 is 8.96 %